# Diffusion-based graph generative methods

### v1

@@ III-B2 Human motion prediction
III-C Other applications
@@
Apart from graph generation applied to the research fields mentioned above, there are also several interesting studies on other applications. SaGess [89] augments DiGress [35] with a generalized divide-and-conquer framework. For 2D graph generation, EDP-GNN [95] is the pioneer in incorporating Score SDEs to discrete graph adjacency matrix generation. Based on this, GraphGDP [36] proposes the denoising kernel named position-enhanced graph score network to make further use of positional information. Instead of sampling only on edges attributes, GSDM [38] leverages Score SDEs on node representations and graph spectrum space to generate graphs with great efficiency, which is shown in results on generic datasets and molecule datasets. Inspired by VAEs, NVDiff [37] takes the VGAE structure and by first samples latent representations of graphs and then decodes them into representations node and edge. SLD [97] takes graph generation task in a similar manner. GraphARM [119] combines autoregressive models with diffusion models by introducing a node-absorbing diffusion process. DruM [96] proposes to directly predict the the destination of the generative process using a mixture of endpoint-conditioned process. EDGE [87] introduces diffusion models to large graph generation. The diffusion process involves gradually remove edges until the graph is empty. The model also manage to avoid the generation of excessive edges by focusing on a small portion of nodes. Specifically designed for hyperbolic graph, HGDM [102] shows promising performance in extracting complex geometric features of hyperbolic embeddings.
@@
In order to make graph-structure prediction in supervised learning, DPM-GSP [84] performs diffusing and denoising process on targets of node classification tasks and achieves outstanding performance. Spatio-temporal graph (STG) forecasting is the problem that requires accurate predictions of signals generated by graphs and their historical observations. DiffSTG [85] is the first research that combines DDPMs with spatio-temporal graph neural networks UGNet. The proposed framework is capable to capture temporal dependencies and spatial correlations. In order to solve NP-complete (NPC) problems, which are fundamental challenges in combinatorial opimization problems, DIFUSCO [91] employs graph-based diffusion models to generate high-quality resolutions. Continuous Gaussian noise and discrete Bernoulli noise are tested and the latter one exhibits better accuracy and scalability. DDM [120] addressses the anisotropic structures in graph data and proposes a novel diffusion model by introducing a data-dependent noise schedule. DiffusionNAG [103] is a novel transferable task-guided framework for neural architecture generation. The model adoptes gradients of the surrogate model to facilitate the model to generate ideal results for given tasks.
@@
Lee et al. utilize diffusion moedels for microstructure reconstruction. Diff-POI [104] samples user spatial preference for next point-of-interest recommendations. Brain Diffuser [121] shapes the structural brain networks from tensor images and shows better ability in exploiting structutal connectivity features and disease-related information. Inspired by nonequilibrium thermodynamics, Lu et al. [122] introduces analog and discrete diffusion with neighbour reprentations to de novo bio-inspired spider web mimics. HouseDiffusion [86] introduces diffusion to vector floorplan generation. Given input graphs where each vertice or edge represents a room or a door, the model samples polygonal loops of vertices and edges. Each loop is composed of a set of sequence of 2D coordinates of corners. Results indicate that HouseDiffusion is able to generate floorplan with great precision. NAP [90] applies DDPM to 3D articulated object synthesis. Articulation tree parameterization is introduced map dynamic articulated objects to articulated graph. By sampling on node and edge attributes, 3D articulated objects are geneated. Illustrations of HouseDiffusion and NAP are shown in Figure 4.
@@ IV-A Evalutaion metrics
IV-A1 Molecule design metrics
@@
IV-A2 Motion generation metrics
@@
Displacement error (DE): the displacement error of all joints in each frame. There are variation such as average displacement error (ADE), final displacement error (FDE), minimen displacement error (MDE), standatd deviation of displacement error (SDE), average final displacement error (AFDE), minimum final displacement error (MFDE), standard deviation of final displacement error (SFDE), Multi-Modal-ADE (MMADE), and Multi-Modal-FDE (MMFDE).
@@ IV-B Datasets
IV-B1 Molecule generation datasets
ZINC [124] is an excellent virtual screening compound library that contains over 750 million purchasable compounds, 230 million compounds of which are in 3D format. GEOM-QM9 [125] and GEOM-Drugs [126] are the most commonly used datasets for 3D molecule and conformation generation. The former one contains molecule conformations in equilibrium state, and is a widely used benchmark dataset for property predictions of molecules in equilibrium state. It contains approximately 133000 molecules, with up to 9 heavy atoms or 29 atoms with hydrogens included.There are only 5 types of atoms: H, C, O, N, and F. GEOM-Drugs [126] includes 37 million molecular conformations for more than 450000 molecules. The maximum number of atoms in this dataset is 181 while the averge is 44. CrossDocked2020 [127] is a dataset of protein-ligand complexes, and known for the inclusion of complexes with both high and low affinity. It contains over 22 million protein-ligand crystal structures. BioLiP [128] is a database of protein-ligand interaction collected from multiple sources get updated weekly. Till June 2023, BioLiP includes over 800000 complexes, over 450000 ligands, and over 400000 proteins. PDBbind [129] is comprised of protein-ligand complexes from Protein Data Bank (PDB). The latest version of it contains over 19000 protein ligand structures. SAbDab is a database composed of all the antibody structures available in the PDB, the total number of antibody structures in SAbDab in the latest version is 7324. For protein generation tasks, PDB provides access to 3D structures of proteins, nucleic acids, and ligand-protein complexes. The most frequently used datasets for molecule generation are summarized in Table II.
@@ IV-B1 Molecule generation datasets
| ZINC | 250,000 molecules | De nove molecule design |
| GEOM-QM9 | 133K molecules and conformations | De novo molecule design, conformation design |
@@
IV-B2 Motion generation
Human3.6M [130], which is composed of 3.6 million human poses divided into 15 categories, is the largest benchmark dataset for human motion studies. Originally, human skeletons in it has 32 joints, but they are not completely adopted in experiments. AMASS [131] is a combination of 18 datasets, which makes it a large one contains 13944 motion sequences. HumanAct12 [132] is a dataset derived from PHSPD [133, 134]. It has 1191 sequences and 90099 frames splitted into 12 categories in total. By re-annotating motions in AMASS and HumanAct12, HumanML3D [135] is a motion-text pair dataset consisting of 14616 sequences with 44970 descriptions. KIT [136] is also a dataset with natural language annotations. Based on AMASS, BABEL [137] contains 13220 sequences seperated into 66289 segments. The dataset is organized into 252 subjects and annotated with 63353 frame-level labels. Being the first 3D motion dataset in the wild, 3DPW [138] has 60 sequences, and 27 two-person sequences annotated with SMPL joints. UESTC [139] contains 24K sequences with 40 subjects. NTU RGB+D [140] is a widely used dataset that is often adopted in the form of NTU RGB+D 60 and NTU RGB+D 120, depending on the number of subjects included. AIST++ [141] is a dataset with high-quality dance motions with 10 different dance motion genres. TSG [142] contains 4 hours of full body motion with speech audio, while ZeroEGGS [143] is consist of two hours of motion and audio data in 19 styles. These two datasets are usually adopted in gestrue genration tasks. Details of the datasets adopted in diffusion-based motion generation research are listed in Table III.
| Dataset | Description | Task |
@@ V Future directions
Combination with AIGC: Recent advances in artificial intelligence generated content (AIGC) arouse widespread research attention and enthusiasm. The pre-training and fine tuning structure of GPT [144, 145] displays promising abilities to generate high-quality samples that are tough for human to distinguish. Combining diffusion-based models with AIGC may result in even better performances.

### v2

@@ III-B2 Human motion prediction
III-B3 Scene Graph Generation
Scene graph generation aim at providing controllable scene synthesis for industrial applications. Specifically, objects are represented by nodes while relationships among them are represented by edges.
Different from previous retrieval-based and semi-generative methods, CommonScenes [104] introduces fully generative model for scene synthesis. It shows the capablity to jointly capture global inter-object relationships and local shape cues. DiffuScene [105] introduces shape latent diffusion for shape retrieval, which performs diffusion process on object attributes and geometries. DiffuseSG [132] proposes joint scene graph imange generation, a novel task in this field. Rather than simply models object attributes, it jointly samples adjacency matrix, node attributes, bounding boxes, and edge attributes. Also focusing on joint scene graph image generation, Wei et al. [107] introduces LLM to facilitate better understanding on graphs and inter-object relationships. To address the fact that methods and data within this field being primarily focused on single scenes, resulting in a lack of flexibility and controllability, InstructScene [106] integrates semantic graph prior and layout decoders to overcome these limitations. It shows comparable performance in multiple downstream tasks in a zero-shot manner.
III-C Generic Graph Generation
In this sub-section, we aim to provide a comprehensive overview on methods that primarily on Generic graph generation. This methods are designed to better adapt diffusion methods to graph generative task without domain restrictions.
@@ III-C Generic Graph Generation
EDGE [52] introduces diffusion models to large graph generation. The diffusion process involves gradually remove edges until the graph is empty. The model also manage to avoid the generation of excessive edges by focusing on a small portion of nodes. Additionally, EDGE++ [111] proposes a degree-specific noise schedule to fully utilize the computational power of diffusion model. The volume-preserved sampling strategy is also introduced to mitigate the inconsistency between behavior of active nodes in diffusion and denoising process. SaGess [114] augments DiGress [35] with a generalized divide-and-conquer framework. For 2D graph generation, EDP-GNN [121] is the pioneer in incorporating Score SDEs to discrete graph adjacency matrix generation. Based on this, GraphGDP [36] proposes the denoising kernel named position-enhanced graph score network to make further use of positional information. Instead of sampling only on edges attributes, GSDM [38] leverages Score SDEs on node representations and graph spectrum space to generate graphs with great efficiency, which is shown in results on generic datasets and molecule datasets. Inspired by VAEs, NVDiff [37] takes the VGAE structure and by first samples latent representations of graphs and then decodes them into representations node and edge. SLD [123] takes graph generation task in a similar manner. GraphARM [150] combines autoregressive models with diffusion models by introducing a node-absorbing diffusion process. Inspired by latent diffusion [151], Graphusion [57] encodes graph into topology-injected latent space, and guides the denoising process with latent self-guidance, which predicts latent variables with pseudo-labels from K-means clustering. In order to preserve invariance, LDM-3DG [58] leverages a pretrained autoencoder to map graph into low-dimension latent space in a data-driven manner. DruM [122] proposes to directly predict the the destination of the generative process using a mixture of endpoint-conditioned process. Specifically designed for hyperbolic graph, HGDM [56] shows promising performance in extracting complex geometric features of hyperbolic embeddings. HypDiff [51] also addresses anisotropy of non-Euclidean structure for graph latent diffusion by introducing two geometric constraints. To generate graph with large scale efficiently, Bergmeister et al. [116] proposes a method that first construct global structure then refines local details.
@@
In order to make graph-structure prediction in supervised learning, DPM-GSP [108] performs diffusing and denoising process on targets of node classification tasks and achieves outstanding performance. Spatio-temporal graph (STG) forecasting is the problem that requires accurate predictions of signals generated by graphs and their historical observations. DiffSTG [109] is the first research that combines DDPMs with spatio-temporal graph neural networks UGNet. The proposed framework is capable to capture temporal dependencies and spatial correlations. DDM [55] addressses the anisotropic structures in graph data and proposes a novel diffusion model by introducing a data-dependent noise schedule. DiffusionNAG [133] is a novel transferable task-guided framework for neural architecture generation. The model adoptes gradients of the surrogate model to facilitate the model to generate ideal results for given tasks.
In the field of knowledge graph embedding (KGE), KGDM [152] transforms the problem of entity prediction into conditional entity generation. On this basis, FDM [153] argues that patterns among entities and ralations of fact come from low-dimensional fact space and introduces a learnable fact embedding model to bridge the gap between diffusion and discrete knowledge graph.
III-D Other Applications
Apart from graph generation applied to the research fields mentioned above, there are also several interesting studies on other applications.
@@ III-D Other Applications
In order to solve NP-complete (NPC) problems, which are fundamental challenges in combinatorial opimization problems, DIFUSCO [115] employs graph-based diffusion models to generate high-quality resolutions. Continuous Gaussian noise and discrete Bernoulli noise are tested and the latter one exhibits better accuracy and scalability. Diff-POI [134] samples user spatial preference for next point-of-interest recommendations. DiffKG [154] integrates diffusion with knowledge graph learning in knowledge-aware recommendations. Inspired by nonequilibrium thermodynamics, Lu et al. [155] introduces analog and discrete diffusion with neighbour reprentations to de novo bio-inspired spider web mimics. HouseDiffusion [110] introduces diffusion to vector floorplan generation. Given input graphs where each vertice or edge represents a room or a door, the model samples polygonal loops of vertices and edges. Each loop is composed of a set of sequence of 2D coordinates of corners. Results indicate that HouseDiffusion is able to generate floorplan with great precision. NAP [113] applies DDPM to 3D articulated object synthesis. Articulation tree parameterization is introduced map dynamic articulated objects to articulated graph. By sampling on node and edge attributes, 3D articulated objects are geneated. Illustrations of HouseDiffusion and NAP are shown in Figure 4.
@@ IV-A Evalutaion metrics
IV-A1 Metrics for AI for scientific discovery tasks
@@
IV-A2 Metrics for computer vision tasks
@@
Displacement error (DE): the displacement error of all joints in each frame. There are variation such as average displacement error (ADE), final displacement error (FDE), minimum displacement error (MDE), standard deviation of displacement error (SDE), average final displacement error (AFDE), minimum final displacement error (MFDE), standard deviation of final displacement error (SFDE), Multi-Modal-ADE (MMADE), and Multi-Modal-FDE (MMFDE).
@@
(3) Scene graph generation
Performance of scene graph generation methods are evaluated from three aspacts, including scene-level fidelity, scene graph consistency, and object-level fidelity. FID and KID are used to evaluate scene-level fidelity. To evaluate the scene graph consistency, the accuracy of a set of relations between generated layout and scene graph is adopted. In terms of object level fidelity, Minimum Matching Distance (MMD), Coverage (COV), and 1-Nearest Neighbor Accuracy (1-NNA) are adopted to compute 3D point clouds between generated samples and groud-truth shapes.
IV-A3 Generic graph generation metrics
In this subsection, we elaborate on metrics for generic graph generation methods.
(1) Structural evaluation metrics
These metrics are used to assess the structural properties of generated graphs and their alignment with original graphs.
Degree Distribution: Measures the distribution of node degrees in the generated graph, reflecting local connectivity patterns.
Clustering Coefficient Distribution: Measures the distribution of clustering coefficients among nodes in the generated graph, indicating the density of connections among neighbors and clustering tendencies.
Orbit Count Distribution: Measures the occurrence frequency of different orbits in the generated graph, typically focusing on orbits with four nodes.
Laplacian Spectrum: Describes the distribution of eigenvalues of the Laplacian matrix of the generated graph.
Characteristic Path Length (CPL): Measures the average shortest path length between nodes in the generated graph.
Assortativity Coefficient (AC): Measures whether nodes with similar degrees in the generated graph tend to be connected to each other.
Triangle Count: Counts the number of triangles (closed loops of three nodes) in the graph.
Square Count: Counts the number of squares (closed loops of four nodes) in the graph.
(2) Quality evaluation metrics
These metrics evaluate the quality of generated graphs and their similarity to the distribution of original data.
Frechet ChemNet Distance (FCD): Evaluates the distance between generated and training graphs in chemical space using activations from the penultimate layer of ChemNet. It assesses the diversity and quality of generated molecules, capturing the similarity in chemical properties between real and generated molecules.
where and are the mean vectors of the generated and training molecules, respectively, and and are their covariance matrices. Tr denotes the trace of a matrix.
Neighborhood Subgraph Pairwise Distance Kernel (NSPDK) MMD: Computes the Maximum Mean Discrepancy (MMD) between generated and test graphs based on node and edge features (degree distribution, clustering coefficient, orbit count distribution, Laplacian spectrum).
where is the sample set of generated graphs, is the sample set of test graphs, and is a kernel function, commonly a Gaussian kernel.
Generation Time: Measures the time required to generate a specified number of graphs.
@@ IV-B Datasets
IV-B1 Datasets for researh on AI for scientific discovery
ZINC [157] is an excellent virtual screening compound library that contains over 750 million purchasable compounds, 230 million compounds of which are in 3D format. GEOM-QM9 [158] and GEOM-Drugs [159] are the most commonly used datasets for 3D molecule and conformation generation. The former one contains molecule conformations in equilibrium state, and is a widely used benchmark dataset for property predictions of molecules in equilibrium state. It contains approximately 133000 molecules, with up to 9 heavy atoms or 29 atoms with hydrogens included.There are only 5 types of atoms: H, C, O, N, and F. GEOM-Drugs [159] includes 37 million molecular conformations for more than 450000 molecules. The maximum number of atoms in this dataset is 181 while the averge is 44. CrossDocked2020 [160] is a dataset of protein-ligand complexes, and known for the inclusion of complexes with both high and low affinity. It contains over 22 million protein-ligand crystal structures. BioLiP [161] is a database of protein-ligand interaction collected from multiple sources get updated weekly. Till June 2023, BioLiP includes over 800000 complexes, over 450000 ligands, and over 400000 proteins. PDBbind [162] is composed of protein-ligand complexes from Protein Data Bank (PDB). The latest version of it contains over 19000 protein ligand structures. SAbDab is a database composed of all the antibody structures available in the PDB, the total number of antibody structures in SAbDab in the latest version is 7324. For protein generation tasks, PDB provides access to 3D structures of proteins, nucleic acids, and ligand-protein complexes. Perov-5 [163, 164] contains perovskite structures with different compositions. Carbon-24 is a dataset of carbon materials simulated 10GPa and MP-20 [165] is a collection of experimentally-generated materials from Material Projects. The most frequently used datasets for molecule generation are summarized in Table LABEL:tab:data_molgen.
@@ IV-B1 Datasets for researh on AI for scientific discovery
ZINC | 250,000 molecules | De nove molecule design, generic graph generation |
GEOM-QM9 | 133K molecules and conformations | De novo molecule design, conformation design, generic graph generation |
@@
Perov-5 | 19K perovskite materials | Crystal structure prediction |
Carbon-24 | 10K carbon materials | Crystal structure prediction |
MP-20 | 45K inorganic materials | Crystal structure prediction |
Community | 500 community graphs | Generic graph generation |
Enzymes | 578 protein graphs | Generic graph generation |
Grid | 100 2D grid graphs | Generic graph generation |
@@
3D-FRONT | 6.8K CAD houses, 18.7K furnished rooms, 7.3K furniture objects | Scene graph generation |
SG-FRONT | 5,7K indoor scenes from 3 scene types with 45K objects | Scene graph generation |
IV-B2 Datasets for research on computer vison
Human3.6M [166], which is composed of 3.6 million human poses divided into 15 categories, is the largest benchmark dataset for human motion studies. Originally, human skeletons in it has 32 joints, but they are not completely adopted in experiments. AMASS [167] is a combination of 18 datasets, which makes it a large one contains 13944 motion sequences. HumanAct12 [168] is a dataset derived from PHSPD [169, 170]. It has 1191 sequences and 90099 frames splitted into 12 categories in total. By re-annotating motions in AMASS and HumanAct12, HumanML3D [171] is a motion-text pair dataset consisting of 14616 sequences with 44970 descriptions. KIT [172] is also a dataset with natural language annotations. Based on AMASS, BABEL [173] contains 13220 sequences seperated into 66289 segments. The dataset is organized into 252 subjects and annotated with 63353 frame-level labels. Being the first 3D motion dataset in the wild, 3DPW [174] has 60 sequences, and 27 two-person sequences annotated with SMPL joints. UESTC [175] contains 24K sequences with 40 subjects. NTU RGB+D [176] is a widely used dataset that is often adopted in the form of NTU RGB+D 60 and NTU RGB+D 120, depending on the number of subjects included. AIST++ [177] is a dataset with high-quality dance motions with 10 different dance motion genres. TSG [178] contains 4 hours of full body motion with speech audio, while ZeroEGGS [179] is consist of two hours of motion and audio data in 19 styles. These two datasets are usually adopted in gestrue genration tasks. 3D-FRONT [180] is a collection of synthetic indoor scene furnished with high-quality objects without detailed discription. Based on 3D-FRONT, SG-FRONT [104] provides annotation labels on a number of object relationships. Details of the datasets are listed in Table II.
@@ V Future directions
Combination with AIGC: Recent advances in artificial intelligence generated content (AIGC) [151, 181] arouse widespread research attention and enthusiasm. The pre-training and fine tuning structure of GPT [182, 183] displays promising abilities to generate high-quality samples that are tough for human to distinguish. Combining diffusion-based models with AIGC may result in even better performances.
@@ VI-A Conclusions
VI-B Comparison with Other Surveys
In the field of diffusion based graph generative methods, surveys by Zhang et al. [19] and Liu et al. [20] share some similarities with our survey. Both their survey mainly focuses on applications of diffusion-based graph generative methods on AI for science discovery. We further elaborate on the applications on computer vision and provide a brief summary on datasets and metrics. We also provide detailed discussion on research that addresses the anisotropic nature of graphs and enhances adaptability of diffusion on graphs.